# See pages - of Main.pdf

### v1

@@ 1.6 Acoustic propagation losses
$\tilde{k}=\beta+i\alpha,$ (23)
@@
$\varepsilon(x)\propto e^{-\alpha x}.$ (25)
@@
$L_{A}\equiv\frac{1}{\alpha},$ (26)
@@
$Q=\frac{\beta}{2\alpha},$ (27)
Figure 6: Measured acoustic resonance at $3~V/\mu m$ bias. solid line represents a Lorentzian fit.
@@
$L_{A}=\frac{2Q}{\beta}=\frac{Q\,\lambda}{\pi}.$ (28)
In our device shown in Fig. 6 ($f_{0}=504~\mathrm{MHz}$, $\gamma/2\pi=1.136~\mathrm{MHz}$), the loaded quality factor is $Q_{L}=f_{0}/(\gamma/2\pi)\approx 444$. With an external coupling fraction $\gamma_{\mathrm{ext}}/\gamma=0.1447$, we obtain $Q_{i}=Q_{L}/(1-\gamma_{\mathrm{ext}}/\gamma)\approx 519$. For an acoustic wavelength $\lambda=8~\mu\mathrm{m}$, the intrinsic strain decay length is
$L_{A,i}=\frac{Q_{i}\lambda}{\pi}\approx\frac{519\times 8~\mu\mathrm{m}}{\pi}\approx 1.32~\mathrm{mm}.$ (29)
@@ 2.1 Calculating the $V_{\pi}L$
To accurately calculate the $V_{\pi}L$ of the acousto-optic modulator, we measure the electrical output coming from the photodetector using an oscilloscope as shown in Fig. 7. We modulate the ring resonator at the 3-dB point to maximize the modulation depth while driving the RF at the Rayleigh wave resonance frequency with 1 dBm RF power. Using the ring transfer function, we can relate the photodetector output peak to peak voltage ($V_{pp}$) to the wavelength shift ($\Delta\lambda$) caused by the $\Delta n_{\mathrm{eff}}$.
@@
Figure 7: Acousto-optic modulation measurement: (a) racetrack response showing the 8.4 pm wavelength shift. (b) Averaged oscilloscope output voltage.
@@
$P_{\mathrm{device}}=P_{\mathrm{input}}(1-|S_{11}|^{2}),$ (30)
@@
$V_{\pi}L=\frac{FSR\;V_{p}\;L}{\Delta\lambda},$ (31)
@@ 2.2 Photoelasticity
$\beta^{(0)}\left(\omega_{0}\right)L=2\pi m,$ (32)
@@
$\beta^{(1)}\left(\omega_{1}\right)L_{c}+\beta^{(0)}\left(\omega_{1}\right)\left(L-L_{c}\right)=2\pi m,$ (33)
@@
$\left(\beta^{(1)}\left(\omega_{1}\right)-\beta^{(0)}\left(\omega_{1}\right)\right)L_{c}+\left(\beta^{(0)}\left(\omega_{1}\right)-\beta^{(0)}\left(\omega_{0}\right)\right)L=$  $0$ (34)
@@
$\delta n_{\mathrm{eff}}=-n_{g}\frac{\delta\omega}{\omega}\frac{L}{L_{c}},$ (35)
@@
$\delta n_{\mathrm{eff}}=\frac{\lambda\;\delta\lambda}{FSR\times L_{c}}.$ (36)
@@
$\delta n_{\mathrm{eff}}$ can also be inferred from perturbation theory [14]
$\delta n_{\mathrm{eff}}=\frac{n_{\mathrm{eff}}}{2}\frac{\left\langle E\right|\Delta\varepsilon\left|E\right\rangle}{\left\langle E\right|\varepsilon\left|E\right\rangle},$ (37)
@@ 2.2.1 Moving boundary contribution
The perturbation term in Eq. (37) arising from domains with shifting material boundaries was first derived in Ref. [8]. For an interface between two materials with permittivities $\varepsilon_{1}$ and $\varepsilon_{2}$, displaced by a mechanical displacement field $\mathbf{u}$, the corresponding moving‑boundary contribution to the permittivity perturbation can be expressed as
$\left\langle E\right|\Delta\varepsilon\left|E\right\rangle=\int ds\ \hat{n}\cdot\vec{u}\left[(\varepsilon_{1}-\varepsilon_{2})\left|E_{\|}\right|^{2}-(\varepsilon_{1}^{-1}-\varepsilon_{2}^{-1})\left|D_{\perp}\right|^{2}\right],$ (38)
@@ 2.2.2 Photoelastic contribution
$\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{M}}=\frac{\epsilon_{0}n_{\mathrm{M}}^{4}n_{\mathrm{eff}}}{2}\frac{\int dr\left(\begin{array}[]{lll}E_{x}^{*}&E_{y}^{*}&E_{z}^{*}\end{array}\right)\left(\begin{array}[]{lll}dB_{1}&dB_{6}&dB_{5}\\ dB_{6}&dB_{2}&dB_{4}\\ dB_{5}&dB_{4}&dB_{3}\end{array}\right)\left(\begin{array}[]{l}E_{x}\\ E_{y}\\ E_{z}\end{array}\right)}{\int E_{\;}^{\ast}\epsilon E\,dr},$ (39)
@@
$\left(\begin{array}[]{c}dB_{1}\\ dB_{2}\\ dB_{3}\\ dB_{4}\\ dB_{5}\\ dB_{6}\end{array}\right)=\left(\begin{array}[]{cccccc}p_{11}&p_{12}&p_{12}&0&0&0\\ p_{12}&p_{11}&p_{12}&0&0&0\\ p_{12}&p_{12}&p_{11}&0&0&0\\ 0&0&0&p_{44}&0&0\\ 0&0&0&0&p_{44}&0\\ 0&0&0&0&0&p_{44}\end{array}\right)\left(\begin{array}[]{c}s_{1}\\ s_{2}\\ s_{3}\\ s_{4}\\ s_{5}\\ s_{6}\end{array}\right),$ (40)
@@
$\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{M}}=-\frac{\epsilon_{0}n_{\mathrm{M}}^{4}n_{\mathrm{eff}}}{2}\frac{\int dr\ [E_{x}^{*}dB_{1}E_{x}+E_{z}^{*}dB_{3}E_{z}+dB_{5}(E_{x}^{*}E_{z}+E_{x}E_{z}^{*})]}{\int E^{*}\epsilon Edr},$ (42)
which gives us the photoelastic contributions of $\mathrm{CSAR}$ and $\mathrm{SiO}_{2}$. In the absence of reported photoelastic coefficients for CSAR 62, we approximate its stress–optic response using those of PMMA[10], a closely related polymeric e-beam resist, noting that the opto-acoustic coupling is dominated by the waveguide core and that the resist contributes only through a weak evanescent optical overlap ($28\%$).
@@ 2.2.3 Effective photoelastic coefficient of SrTiO3
$\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{STO}}$ $=\frac{\epsilon_{0}n_{\mathrm{STO}}^{4}n_{\mathrm{eff}}}{2}\frac{\int dr\ [(s_{1}+s_{3})(|E_{x}|^{2}+|E_{z}|^{2})+2s_{5}\mathrm{Re}(E_{x}^{*}E_{z})]}{\int E^{*}\epsilon Edr}p_{\mathrm{eff}},$ (43)
@@
$p_{\mathrm{eff}}=\frac{1}{\mathcal{L}}(\delta n_{\mathrm{eff}}-\delta n_{\mathrm{eff},\mathrm{MB}}-\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{CSAR}}-\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{SiO}_{2}}).$ (44)
Table 1 summarizes the analysis parameters and results. Finally, we benchmark the extracted effective photoelastic coefficient of SrTiO3 against that of BaTiO3, using identical simulated strain and optical field distributions. The photoelastic tensor elements for BaTiO3 are taken from literature [22]
$p_{eff,BTO}=\frac{\int dr\left[E_{x}^{*}dB_{1,BTO}E_{x}+E_{y}^{*}dB_{2,BTO}E_{y}+E_{z}^{*}dB_{3,BTO}E_{z}+dB_{5,BTO}\left(E_{x}^{*}E_{z}+E_{x}E_{z}^{*}\right)\right]}{\int dr\left[\left(s_{1}+s_{3}\right)\left(\left|E_{x}\right|^{2}+\left|E_{y}\right|^{2}+\left|E_{z}\right|^{2}\right)+2s_{5}\operatorname{Re}\left(E_{x}^{*}E_{z}\right)\right]},$ (45)
@@
Table 1: Key parameters and extracted quantities used in this work.
@@ 2.3 Optical response with increased RF power
Figure 8: (a) Measured optical transmission spectra of the ring resonator for different RF powers applied to the IDT. (b) Simulated optical transmission.
Figure 8a shows the optical transmission as a function of the applied RF power. In the unresolved-sideband regime $\Omega\ll\kappa$, the optical cavity responds adiabatically to the mechanical motion, such that the optomechanical interaction induces resonance-frequency jitter of the cavity. The measured transmission, which represents a time average over this frequency jitter, therefore exhibits an apparent broadening relative to the intrinsic cavity linewidth. Additionally, we observed that increasing the RF power leads to a slight increase in the sample temperature, accompanied by a blue shift of the optical resonance. Using temperature-dependent measurements of the ring response, we correlated the observed thermal shift with the actual device temperature, estimating a temperature increase of 6.26 K at 10 dBm RF power.
The broadening of the measured optical response under high RF power can be numerically modeled assuming a weak optical input using the equation of the intracavity field amplitude $a(t)$ [14]
$\dot{a}(t)=\left(i\Delta-\frac{\kappa}{2}\right)a(t)-i\,\delta\omega\cos(\Omega t)\,a(t)+\sqrt{\kappa_{e}}\,a_{\mathrm{in}},$ (46)
@@
$\left(\frac{\kappa}{2}-i(\Delta+n\Omega)\right)a_{n}+i\frac{\delta\omega}{2}\left(a_{n-1}+a_{n+1}\right)=\sqrt{\kappa_{e}}\,a_{\mathrm{in}}\,\delta_{n0}.$ (48)
@@
$a_{\mathrm{out}}(t)=a_{\mathrm{in}}-\sqrt{\kappa_{e}}\,a(t),$ (49)
@@
$T_{\mathrm{DC}}=\left\langle\frac{|a_{\mathrm{out}}(t)|^{2}}{|a_{\mathrm{in}}|^{2}}\right\rangle=\frac{\left|a_{\mathrm{in}}-\sqrt{\kappa_{e}}\,a_{0}\right|^{2}+\kappa_{e}\sum_{n\neq 0}|a_{n}|^{2}}{|a_{\mathrm{in}}|^{2}}.$ (50)
The quasi-DC transmission spectrum $T_{\mathrm{DC}}(\lambda)$ is then obtained by evaluating Eq. (50) as a function of laser detuning $\Delta(\lambda)$, as shown in Fig. 8b.
@@ 3 Materials overview
Figure 9: Piezoelectric tensor elements of commonly used materials. Effective parameters in SrTiO3, arising from twin domain boundaries, are denoted by stars. [18, 19, 1, 16, 9, 15].
From an electromechanical perspective, a central figure of merit is the electromechanical coupling rate $g_{\mathrm{EM}}$. In piezoelectric resonators, $g_{\mathrm{EM}}$ scales with the square root of the effective electromechanical coupling coefficient $k^{2}$[20], which is determined by the relevant piezoelectric tensor elements $e_{ij}$, as well as by device geometry and modal overlap. Materials with large intrinsic piezoelectric coefficients, such as AlN and LiNbO3, naturally support large $k^{2}$ and can therefore achieve strong electromechanical coupling in compact geometries. In contrast, GaAs exhibits a comparatively weak piezoelectric response, resulting in smaller $k^{2}$ and correspondingly reduced $g_{\mathrm{EM}}$, unless additional resonant enhancement or impedance-matching strategies are employed. Figure 9 compares the dominant piezoelectric tensor elements for the materials considered here.
At cryogenic temperatures, the piezoelectric response of ferroelectric materials generally decreases as the operating temperature moves further away from the Curie point. The magnitude of this reduction is material dependent; for example, LiNbO3 exhibits a relatively modest decrease of approximately 6% at 1.3 K [4].
Optomechanical performance, by contrast, is primarily characterized by the single-photon optomechanical coupling rate $g_{0}$, which depends on the photoelastic tensor, the refractive index, and the achievable optical mode confinement. While conventional piezoelectric materials provide moderate photoelastic coupling, hybrid photonic platforms enable access to materials with substantially larger photoelastic coefficients. In particular, SiN-on-SrTiO3 and all-SrTiO3 waveguides would leverage a strong photoelastic response enabling enhanced optomechanical interaction strengths. Table 2 summarizes the relevant photoelastic tensor elements for representative optical guiding materials.
Table 2: Photoelastic tensor elements of commonly used optical guiding materials, where bold numbers indicate the high values [7, 17, 24, 22].

### v2

@@ 1.6 Acoustic propagation losses
$\tilde{k}=\beta+i\alpha,$ (S23)
@@
$\varepsilon(x)\propto e^{-\alpha x}.$ (S25)
@@
$L_{A}\equiv\frac{1}{\alpha},$ (S26)
@@
$Q=\frac{\beta}{2\alpha},$ (S27)
[…]
@@
$L_{A}=\frac{2Q}{\beta}=\frac{Q\,\lambda}{\pi}.$ (S28)
In our device shown in Fig. S6 ($f_{0}=504~\mathrm{MHz}$, $\gamma/2\pi=1.136~\mathrm{MHz}$), the loaded quality factor is $Q_{L}=f_{0}/(\gamma/2\pi)\approx 444$. With an external coupling fraction $\gamma_{\mathrm{ext}}/\gamma=0.1447$, we obtain $Q_{i}=Q_{L}/(1-\gamma_{\mathrm{ext}}/\gamma)\approx 519$. For an acoustic wavelength $\lambda=8~\mu\mathrm{m}$, the intrinsic strain decay length is
$L_{A,i}=\frac{Q_{i}\lambda}{\pi}\approx\frac{519\times 8~\mu\mathrm{m}}{\pi}\approx 1.32~\mathrm{mm}.$ (S29)
@@ 2.1 Calculating the $V_{\pi}L$
To accurately calculate the $V_{\pi}L$ of the acousto-optic modulator, we measure the electrical output coming from the photodetector using an oscilloscope as shown in Fig. S7. We modulate the ring resonator at the 3-dB point to maximize the modulation depth while driving the RF at the Rayleigh wave resonance frequency with 1 dBm RF power. Using the ring transfer function, we can relate the photodetector output peak to peak voltage ($V_{pp}$) to the wavelength shift ($\Delta\lambda$) caused by the $\Delta n_{\mathrm{eff}}$.
@@
Fig. S7: Acousto-optic modulation measurement. (a) Racetrack response showing the 8.4 pm wavelength shift. (b) Averaged oscilloscope output voltage.
@@
$P_{\mathrm{device}}=P_{\mathrm{input}}(1-|S_{11}|^{2}),$ (S30)
@@
$V_{\pi}L=\frac{FSR\;V_{p}\;L}{\Delta\lambda},$ (S31)
@@ 2.2 Photoelasticity
$\beta^{(0)}\left(\omega_{0}\right)L=2\pi m,$ (S32)
@@
$\beta^{(1)}\left(\omega_{1}\right)L_{c}+\beta^{(0)}\left(\omega_{1}\right)\left(L-L_{c}\right)=2\pi m,$ (S33)
@@
$\left(\beta^{(1)}\left(\omega_{1}\right)-\beta^{(0)}\left(\omega_{1}\right)\right)L_{c}+\left(\beta^{(0)}\left(\omega_{1}\right)-\beta^{(0)}\left(\omega_{0}\right)\right)L=$  $0$ (S34)
@@
$\delta n_{\mathrm{eff}}=-n_{g}\frac{\delta\omega}{\omega}\frac{L}{L_{c}},$ (S35)
@@
$\delta n_{\mathrm{eff}}=\frac{\lambda\;\delta\lambda}{FSR\times L_{c}}.$ (S36)
@@
$\delta n_{\mathrm{eff}}$ can also be inferred from perturbation theory (?)
$\delta n_{\mathrm{eff}}=\frac{n_{\mathrm{eff}}}{2}\frac{\left\langle E\right|\Delta\varepsilon\left|E\right\rangle}{\left\langle E\right|\varepsilon\left|E\right\rangle},$ (S37)
@@ 2.2.1 Moving boundary contribution
The perturbation term in Eq. (S37) arising from domains with shifting material boundaries was first derived in Ref.  (?). For an interface between two materials with permittivities $\varepsilon_{1}$ and $\varepsilon_{2}$, displaced by a mechanical displacement field $\mathbf{u}$, the corresponding moving‑boundary contribution to the permittivity perturbation can be expressed as
$\left\langle E\right|\Delta\varepsilon\left|E\right\rangle=\int ds\ \hat{n}\cdot\vec{u}\left[(\varepsilon_{1}-\varepsilon_{2})\left|E_{\|}\right|^{2}-(\varepsilon_{1}^{-1}-\varepsilon_{2}^{-1})\left|D_{\perp}\right|^{2}\right],$ (S38)
@@ 2.2.2 Photoelastic contribution
$\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{M}}=\frac{\epsilon_{0}n_{\mathrm{M}}^{4}n_{\mathrm{eff}}}{2}\frac{\int dr\left(\begin{array}[]{lll}E_{x}^{*}&E_{y}^{*}&E_{z}^{*}\end{array}\right)\left(\begin{array}[]{lll}dB_{1}&dB_{6}&dB_{5}\\ dB_{6}&dB_{2}&dB_{4}\\ dB_{5}&dB_{4}&dB_{3}\end{array}\right)\left(\begin{array}[]{l}E_{x}\\ E_{y}\\ E_{z}\end{array}\right)}{\int E_{\;}^{\ast}\epsilon E\,dr},$ (S39)
@@
$\left(\begin{array}[]{c}dB_{1}\\ dB_{2}\\ dB_{3}\\ dB_{4}\\ dB_{5}\\ dB_{6}\end{array}\right)=\left(\begin{array}[]{cccccc}p_{11}&p_{12}&p_{12}&0&0&0\\ p_{12}&p_{11}&p_{12}&0&0&0\\ p_{12}&p_{12}&p_{11}&0&0&0\\ 0&0&0&p_{44}&0&0\\ 0&0&0&0&p_{44}&0\\ 0&0&0&0&0&p_{44}\end{array}\right)\left(\begin{array}[]{c}s_{1}\\ s_{2}\\ s_{3}\\ s_{4}\\ s_{5}\\ s_{6}\end{array}\right),$ (S40)
@@
$\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{M}}=-\frac{\epsilon_{0}n_{\mathrm{M}}^{4}n_{\mathrm{eff}}}{2}\frac{\int dr\ [E_{x}^{*}dB_{1}E_{x}+E_{z}^{*}dB_{3}E_{z}+dB_{5}(E_{x}^{*}E_{z}+E_{x}E_{z}^{*})]}{\int E^{*}\epsilon Edr},$ (S42)
which gives us the photoelastic contributions of $\mathrm{CSAR}$ and $\mathrm{SiO}_{2}$. In the absence of reported photoelastic coefficients for CSAR 62, we approximate its stress–optic response using those of PMMA (?), a closely related polymeric e-beam resist, noting that the opto-acoustic coupling is dominated by the waveguide core and that the resist contributes only through a weak evanescent optical overlap ($28\%$).
@@ 2.2.3 Effective photoelastic coefficient of SrTiO3
$\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{STO}}$ $=\frac{\epsilon_{0}n_{\mathrm{STO}}^{4}n_{\mathrm{eff}}}{2}\frac{\int dr\ [(s_{1}+s_{3})(|E_{x}|^{2}+|E_{z}|^{2})+2s_{5}\mathrm{Re}(E_{x}^{*}E_{z})]}{\int E^{*}\epsilon Edr}p_{\mathrm{eff}},$ (S43)
@@
$p_{\mathrm{eff}}=\frac{1}{\mathcal{L}}(\delta n_{\mathrm{eff}}-\delta n_{\mathrm{eff},\mathrm{MB}}-\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{CSAR}}-\delta n_{\mathrm{eff},\mathrm{PE},\mathrm{SiO}_{2}}).$ (S44)
Table S1 summarizes the analysis parameters and results. Finally, we benchmark the extracted effective photoelastic coefficient of SrTiO3 against that of BaTiO3, using identical simulated strain and optical field distributions. The photoelastic tensor elements for BaTiO3 are taken from literature  (?)
$p_{\mathrm{eff,BTO}}=\frac{\int dr\left[E_{x}^{*}dB_{1,BTO}E_{x}+E_{y}^{*}dB_{2,BTO}E_{y}+E_{z}^{*}dB_{3,BTO}E_{z}+dB_{5,BTO}\left(E_{x}^{*}E_{z}+E_{x}E_{z}^{*}\right)\right]}{\int dr\left[\left(s_{1}+s_{3}\right)\left(\left|E_{x}\right|^{2}+\left|E_{y}\right|^{2}+\left|E_{z}\right|^{2}\right)+2s_{5}\operatorname{Re}\left(E_{x}^{*}E_{z}\right)\right]},$ (S45)
@@
Table S1: Key parameters and extracted quantities used in this work.
@@ 2.3 Optical response with increased RF power
Fig. S8: Optical response vs. RF power. (a) Measured optical transmission spectra of the ring resonator for different RF powers applied to the IDT. (b) Simulated optical transmission.
Fig. S8a shows the optical transmission as a function of the applied RF power. In the unresolved-sideband regime $\Omega\ll\kappa$, the optical cavity responds adiabatically to the mechanical motion, such that the optomechanical interaction induces resonance-frequency jitter of the cavity. The measured transmission, which represents a time average over this frequency jitter, therefore exhibits an apparent broadening relative to the intrinsic cavity linewidth. Additionally, we observed that increasing the RF power leads to a slight increase in the sample temperature, accompanied by a blue shift of the optical resonance. Using temperature-dependent measurements of the ring response, we correlated the observed thermal shift with the actual device temperature, estimating a temperature increase of 6.26 K at 10 dBm RF power.
The broadening of the measured optical response under high RF power can be numerically modeled assuming a weak optical input using the equation of the intracavity field amplitude $a(t)$ (?)
$\dot{a}(t)=\left(i\Delta-\frac{\kappa}{2}\right)a(t)-i\,\delta\omega\cos(\Omega t)\,a(t)+\sqrt{\kappa_{e}}\,a_{\mathrm{in}},$ (S46)
@@
$\left(\frac{\kappa}{2}-i(\Delta+n\Omega)\right)a_{n}+i\frac{\delta\omega}{2}\left(a_{n-1}+a_{n+1}\right)=\sqrt{\kappa_{e}}\,a_{\mathrm{in}}\,\delta_{n0}.$ (S48)
@@
$a_{\mathrm{out}}(t)=a_{\mathrm{in}}-\sqrt{\kappa_{e}}\,a(t),$ (S49)
@@
$T_{\mathrm{DC}}=\left\langle\frac{|a_{\mathrm{out}}(t)|^{2}}{|a_{\mathrm{in}}|^{2}}\right\rangle=\frac{\left|a_{\mathrm{in}}-\sqrt{\kappa_{e}}\,a_{0}\right|^{2}+\kappa_{e}\sum_{n\neq 0}|a_{n}|^{2}}{|a_{\mathrm{in}}|^{2}}.$ (S50)
The quasi-DC transmission spectrum $T_{\mathrm{DC}}(\lambda)$ is then obtained by evaluating Eq. (S50) as a function of laser detuning $\Delta(\lambda)$, as shown in Fig. S8b.
@@ 3 Materials overview
Fig. S9: Piezoelectric tensor comparison. Piezoelectric tensor elements of commonly used materials. Effective parameters in SrTiO3, arising from twin domain boundaries, are denoted by stars. (?, ?, ?, ?, ?, ?).
From an electromechanical perspective, a central figure of merit is the electromechanical coupling rate $g_{\mathrm{EM}}$. In piezoelectric resonators, $g_{\mathrm{EM}}$ scales with the square root of the effective electromechanical coupling coefficient $k^{2}$ (?), which is determined by the relevant piezoelectric tensor elements $e_{ij}$, as well as by device geometry and modal overlap. Materials with large intrinsic piezoelectric coefficients, such as AlN and LiNbO3, naturally support large $k^{2}$ and can therefore achieve strong electromechanical coupling in compact geometries. In contrast, GaAs exhibits a comparatively weak piezoelectric response, resulting in smaller $k^{2}$ and correspondingly reduced $g_{\mathrm{EM}}$, unless additional resonant enhancement or impedance-matching strategies are employed. Fig. S9 compares the dominant piezoelectric tensor elements for the materials considered here.
At cryogenic temperatures, the piezoelectric response of ferroelectric materials generally decreases as the operating temperature moves further away from the Curie point. The magnitude of this reduction is material dependent; for example, LiNbO3 exhibits a relatively modest decrease of approximately 6% at 1.3 K (?).
Optomechanical performance, by contrast, is primarily characterized by the single-photon optomechanical coupling rate $g_{0}$, which depends on the photoelastic tensor, the refractive index, and the achievable optical mode confinement. While conventional piezoelectric materials provide moderate photoelastic coupling, hybrid photonic platforms enable access to materials with substantially larger photoelastic coefficients. In particular, SiN-on-SrTiO3 and all-SrTiO3 waveguides would leverage a strong photoelastic response enabling enhanced optomechanical interaction strengths. Table S2 summarizes the relevant photoelastic tensor elements for representative optical guiding materials.
Table S2: Photoelastic tensor elements of commonly used optical guiding materials, where bold numbers indicate the high values (?, ?, ?, ?).